# Mirror: A Multiple-perspective Self-Reflection Method for Knowledge-rich Reasoning

### v1

@@ 2 Related Work
Extensive research (Honovich et al., 2022; Xie et al., 2023) has been conducted to enhance LLMs through the concept of self-reflection, where LLMs learn from automatically generated feedback to understand and reflect on their own outputs. This feedback can stem from various sources: the LLM itself (Madaan et al., 2023; Shinn et al., 2023), a separately trained critic module (Gou et al., 2023b; Peng et al., 2023) or external sources (Yao et al., 2023b), such as Wikipedia or an Internet Browser. Gou et al. (2023b); Peng et al. (2023) argued that evaluators trained on task-oriented feedback offer superior performance. For example, Refiner Paul et al. (2023) took context and hypotheses as input to generate templates-based feedback for various error types. Recent studies (Peng et al., 2023; Shinn et al., 2023; Hao et al., 2023) have fully utilized the in-context learning capability of LLMs, prompting them to generate high-quality feedback based on their previous generation or potential templates. Madaan et al. (2023) proposed multiple task-oriented metrics and prompted LLMs to evaluate their own outputs based on these criteria. Similarly, Peng et al. (2023); Glaese et al. (2022) adopted external tools to predict multi-facet human preference scores. Our solution aligns with this trend by aiming to provide informative and customized instructions tailored to the specific task and query. Moreover, it seeks to achieve this without relying on human intervention or external tools, thereby rendering self-refinement more feasible in practice.
@@
Recently, tree-based reasoning has attracted significant attention, such as Tree-of-Thought (ToT) (Yao et al., 2023a), Grace (Khalifa et al., 2023), and SelfEval-Decoding (Xie et al., 2023). At each reasoning step, ToT adopts breadth-first search and depth-first search, while the latter two methods select the top-$k$ scoring candidates during the decoding process. Moreover, Monte-Carlo Tree Search (MCTS) is one of the popular search algorithms (Swiechowski et al., 2023), which strikes a balance between exploitation and exploration. Some existing approaches establish a reinforcement learning framework to maximize reward through learning optimal actions/states (Du et al., 2023a; Parthasarathy et al., 2023; Zhu et al., 2023). Other studies fully utilize the capability of LLMs for interaction and feedback generation. For instance, RAP (Hao et al., 2023) leveraged step-wise rewards from interactions with the world model to decompose and solve the problem step-by -step, rather than a iterative manner. LATS (Zhou et al., 2023) was the first work in leveraging MCTS for self-reflection. However, their feedback contains information from comparisons with ground truth, which is not applicable in our case. Instead, our approach, Mirror has no access to gold labels, and we incorporate a novel diversity reward to avoid the inefficient search in the reflection iteration.
@@ 4.1 Problem Setup
Given a question, the Reasoner is to arrive at the final answer through interacting with a Navigator. We consider a Markov Decision Process (MDP) defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{\pi},\gamma,\mathcal{R})$, where the $s_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$ denote the state and action, respectively in the $t$-th reflection iteration. In the context of multiple-choice question, $a_{t}$ is the direction generated by the Navigator, and $s_{t}$ is the response generated by the Reasoner, including the answer to the question and the rationale behind. $\mathcal{R}(s,a)$ is the reward function. Therefore, we have state transition distribution $\mathcal{P}(s_{t}|s_{t-1},a_{t-1})$ and action generation distribution $\mathcal{\pi}(a_{t}|s_{t},q,p_{0},\mathcal{R})$, where $p_{0}$ is the prompt for the Navigator to generate direction $a_{t}$. It is nontrivial to obtain frequent rewards that incentivize self-refinement progress without access to the ground truth. Therefore, we turn to an intrinsically motivated planning algorithm, i.e., Monte-Carlo Tree Search (MCTS) (Kocsis and Szepesvári, 2006; Browne et al., 2012; Swiechowski et al., 2023) to efficiently explore the environment augmenting rewards with auxiliary objectives (Mu et al., 2022; Du et al., 2023b).
@@ 4.2 Multiple-perspective Assessment
Motivated by the empirical results in § 3.1 regarding knowledge-grounding, we propose to employ an advanced consistency-based method as a surrogate for factual correctness when external resources are unavailable. This method considers both intra- and inter-consistency of the generated responses. Specially, we employ the Navigator for $K$ question-oriented direction generation, $a\sim\pi(a_{t}|q,s_{t},p_{0},\mathcal{R})$. These $K$ directions are intended to provide diverse perspectives for problem-solving, with the agreement among guided responses representing inter-consistency. Meanwhile, the confidence in self-consistency (Wang et al., 2023) serves as the measure of intra-consistency. To integrate consistency considerations into the assessment per reflection iteration, we use intra-consistency to determine whether the Reasoner should accept its initial response. If the intra-consistency surpasses a threshold $T_{0}$, we consider it as the final result; otherwise, we integrate the inter-consistency as an indicator for stopping criteria in subsequent reflection iterations. We derive the final answer when the inter-consistency exceeds $T_{0}$ or when reach the predefined maximum iterations, selecting the final answer with the highest consistency score. This inter-consistency also becomes part of reward $\mathcal{R}_{\text{consistency}}$ for the current state and contribute to the direction generation. We compare with majority voting in Table 4 to illustrate the efficiency of our assessment strategy.
@@ 4.3 Diverse and Valid Search Space
Obtaining a meaningful and diverse action space is challenging due to the absence of a dense and well-defined reward function in the planning algorithm. One of the predominant reasons is that different action sequences can lead to similar outcomes (Baranes and Oudeyer, 2013). In our context, considering the limitation of LLMs in following instructions, the Reasoner may ignore the differences among multiple directions and generate identical responses merely based on the question. Therefore, some intrinsically motivated reinforcement learning algorithms choose to explore outcomes rather than actions (Oudeyer and Kaplan, 2007; Ladosz et al., 2022). MCTS addresses the limitation of sparse rewards by visiting novel states or transitions through random exploration (Du et al., 2023b). The most popular algorithm in the MCTS family, Upper Confidence Bound for Trees (UCT) (Kocsis and Szepesvári, 2006) is treated as the choice of child node, $\text{UCT}=\overline{\mathcal{R}}_{j}+2C_{p}\sqrt{\frac{2\,\text{In}\,N(n)}{N(n_{j})}}$, where ¯$\overline{\mathcal{R}}_{j}$ is the average reward for child node $j$, while the second term encourages sampling from nodes whose children are less visited. $N(n)$ is the number of times current node (parent) has been visited in previous iterations, and $N(n_{j})$ is times of the child node has been visited. The $C_{p}>0$ is a constant to control balance between exploitation (first term) and exploration (second term). In our case, we specifically promote diversity between the parent and child node, i.e., the response in previous attempt $s_{t-1}$ and the current attempt $s_{t}$ [^3] We conducted experiments by filtering $K$ distinct directions based on their semantic similarity derived from SentenceBERT (Reimers and Gurevych, 2019), but the performances did not match those achieved by directly constraining the diversity of outcomes.. For multiple-choice questions in MMLU, we assess if the predicted choices are the same across two reflection iterations. The discrepancy in responses indicates the alleviation of null direction space and the avoidance of being stuck, especially given the relatively low consistency with the response from the previous iteration. The relationship between task performance and the diversity of responses in the generated tree, as illustrated in Figure 5, confirms our motivation for diversity enhancement. However, maximizing diversity of outcomes may not always be enough, as less relevant states might be collected (Du et al., 2023b). Therefore, we filter out states whose associated responses are not in the correct form, such as failing to provide a final choice, or refusing to answer questions for moral considerations.
@@ 5 Can Mirror Steer LLMs in Iterative Improvements?
We evaluate our proposed Mirror on MMLU and FEVER (Thorne et al., 2018). FEVER is a fact-checking dataset featuring three labels for knowledge-rich statements, i.e., supports, refutes and not enough info.
@@ 5.1 Experimental Setup and Results
The evaluation models are GPT-3.5, Llama2-13B-Chat (Touvron et al., 2023), and Vicuna-v1.5-7B (Zheng et al., 2023)  [^4] We denote them as Llama13B and Vicuna13B for simplicity. Experiment details can be found in Appendix C.. We equip the LLMs with different reasoning mechanisms, including Chain-of-Thought (CoT) (Wei et al., 2022), Self-consistency (Wang et al., 2023), Self-Correction (Huang et al., 2023) and Reflexion(w.GT) Shinn et al. (2023). We implement CoT by prompting LLMs to first generate step-by-step thoughts and then generate answers based on those thoughts. We repeat this process for five times, resulting in Self-Consistency${}^{(5)}$. The remaining two methods are self-improvement techniques where LLMs are first prompted to generate reflections, followed by updating their current response accordingly if applicable. Self-Correction relies on LLM’s internal knowledge for answer assessment, while Reflexion compares the current answer with the ground truth for evaluation.
Figure 4: Reasoning process of self-correction and Mirror. Text in red are generated directions.
@@
Self-Consistency${}^{(5)}$ 0.67 0.68 0.58 0.64 0.61
Self-Correct${}^{(2)}$ 0.63 0.62 0.55 0.54 0.55
@@
Self-Consistency${}^{(5)}$ 0.45 0.60 0.49 0.57 0.46
Self-Correct${}^{(2)}$ 0.42 0.52 0.53 0.45 0.36
@@
Self-Consistency${}^{(5)}$ 0.50 0.62 0.53 0.60 0.43
Self-Correct${}^{(2)}$ 0.43 0.49 0.42 0.49 0.38
@@ Results.
The results are shown in Table 2. By comparing CoT with Self-Correction, we observe the performance degradation after two rounds of self-Correction across almost all datasets and models. This observation aligns with our findings in §3.1 and in Huang et al. (2023). Equipped with self-consistency${}^{(5)}$, significant performance improvements are evident across all settings. Mirror considers additional inter-consistency, achieves the most notable improvements, with a relative increase of more than 15% across the three models. Figure 4 illustrates the reasoning process of Self-correction and Mirror. Both methods fail to answer correctly in the first trial. With question-oriented direction, the Reasoner better identify errors in the initial response, such as, the error in score direction and inconsistency between rationales and selection. The consistency-based criteria built in the tree further improves the fact assessment. During backpropagation, node $s^{(1)}_{1}$ receives a higher reward, leading to the leftmost reasoning path (details of direction $a^{(1)}_{1},a^{(1)}_{2},a^{(2)}_{2}$ and corresponding responses are shown in the text frame). By contrast, Self-correction seems to engage in groundless inference by switching answers without explicit clues. Even comparing Mirror with Relexion(w.GT), we find comparable results for GPT-3.5 on the STEM dataset, for Llama on all datasets except for STEM and for Vicuna on STEM and Humanity. From the perspective of the model, the average improvements over baselines for GPT-3.5 are particularly prominent, partly explained by its better ability to adhere to provided directions. This can also explain the marginal improvements even ground truth are accessible to the smaller models.
@@ 5.2 Analysis
GPT-3.5:
CoT
0.68 0.58
+ FixedDirect
0.73 0.60
+ GenerativeDirect
0.78 0.64
Llama13B:
CoT
0.46 0.40
+ FixedDirect
0.43 0.39
+ GenerativeDirect
0.49 0.45
Vicuna13B:
CoT
0.48 0.42
+ FixedDirect
0.51 0.43
+ GenerativeDirect
0.55 0.45
@@
Figure 5: The Accuracy (acc) and the percentage of samples where the ground truth is included in the tree (ans-presence), with different sizes of search space (Num). Results for GPT-3.5 and Llama13B are in Figure 3(a) and 3(b).
@@
As discussed in Section 3.1, LLMs struggle to assess the correctness of knowledge-rich statements, a capability that can be consistently enhanced through self-consistency. We further reform the majority-voting assessment process by considering the inter-consistency built in the hierarchical decision-making tree. To study the effects of our answer assessment criteria described in §4.2, we compare them with two other voting methods, i.e., self-consistency (majority vote for 5 CoT-generated reasoning trajectories) and majority vote within our generated tree-trajectories. We average the results from Table 2 for CoT and Self-consistency${}^{(5)}$ across four domains in MMLU and denote them as CoT${}^{(1)}$ and CoT${}^{(5)}$, respectively. For Majority${}^{(tree)}$, we select the final answer through majority-voting among all intermediate nodes in our generated tree-trajectories. The results of different final answer assessments are presented in Table 4. We observe a performance increase after applying majority-voting in the CoT settings, while this simple strategy doesn’t yield improvements in the generated tree. This is because undesirable responses may be generated during the node expanding phase, and majority voting treats all nodes equally. In contrast, our reward-based search tends to focus on reliable nodes with higher confidence in each reflection step, thereby avoiding less desirable nodes.
@@
GPT-3.5: CoT${}^{(1)}$ 0.60 0.58
CoT${}^{(5)}$ 0.64 0.61
+Majority${}^{(tree)}$ 0.69 0.59
+Reward Search${}^{(tree)}$ 0.73 0.64
Llama13B: CoT${}^{(1)}$ 0.49 0.40
CoT${}^{(5)}$ 0.53 0.46
+Majority${}^{(tree)}$ 0.58 0.50
+Reward Search${}^{(tree)}$ 0.60 0.54
Vicuna13B: CoT${}^{(1)}$ 0.51 0.39
CoT${}^{(5)}$ 0.56 0.43
+Majority${}^{(tree)}$ 0.59 0.43
+Reward Search${}^{(tree)}$ 0.60 0.46

### v2

@@ 2 Related Work
Extensive research (Honovich et al., 2022; Xie et al., 2023b) has been conducted to enhance LLMs through the concept of self-reflection, where LLMs learn from automatically generated feedback to understand and reflect on their own outputs. This feedback can stem from various sources: the LLM itself (Madaan et al., 2023; Shinn et al., 2023), a separately trained critic module (Gou et al., 2023b; Peng et al., 2023) or external sources (Yao et al., 2023b), such as Wikipedia or an Internet Browser. Gou et al. (2023b); Peng et al. (2023) argued that evaluators trained on task-oriented feedback offer superior performance. For example, Refiner Paul et al. (2023) took context and hypotheses as input to generate templates-based feedback for various error types. Recent studies (Peng et al., 2023; Shinn et al., 2023; Hao et al., 2023) have fully utilized the in-context learning capability of LLMs, prompting them to generate high-quality feedback based on their previous generation or potential templates. Madaan et al. (2023) proposed multiple task-oriented metrics and prompted LLMs to evaluate their own outputs based on these criteria. Similarly, Peng et al. (2023); Glaese et al. (2022) adopted external tools to predict multi-facet human preference scores. Our solution aligns with this trend by aiming to provide informative and customized instructions tailored to the specific task and query. Moreover, it seeks to achieve this without relying on human intervention or external tools, thereby rendering self-refinement more feasible in practice.
@@
Recently, tree-based reasoning has attracted significant attention, such as Tree-of-Thought (ToT) (Yao et al., 2023a), Grace (Khalifa et al., 2023), and SelfEval-Decoding (Xie et al., 2023b). At each reasoning step, ToT adopts breadth-first search and depth-first search, while the latter two methods select the top-$k$ scoring candidates during the decoding process. Moreover, Monte-Carlo Tree Search (MCTS) is one of the popular search algorithms (Swiechowski et al., 2023), which strikes a balance between exploitation and exploration. Some existing approaches establish a reinforcement learning framework to maximize reward through learning optimal actions/states (Du et al., 2023a; Parthasarathy et al., 2023; Zhu et al., 2023). Other studies fully utilize the capability of LLMs for interaction and feedback generation. For instance, RAP (Hao et al., 2023) leveraged step-wise rewards from interactions with the world model to decompose and solve the problem step-by -step, rather than a iterative manner. LATS (Zhou et al., 2023) was the first work in leveraging MCTS for self-reflection. However, their feedback contains information from comparisons with ground truth, which is not applicable in our case. Instead, our approach, Mirror has no access to gold labels, and we incorporate a novel diversity reward to avoid the inefficient search in the reflection iteration.
@@ 4.1 Problem Setup
Given a question, the Reasoner is to arrive at the final answer through interacting with a Navigator. We consider a Markov Decision Process (MDP) defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{\pi},\gamma,\mathcal{R})$, where the $s_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$ denote the state and action, respectively in the $t$-th reflection iteration. In the context of multiple-choice question, $a_{t}$ is the direction generated by the Navigator, and $s_{t}$ is the response generated by the Reasoner, including the answer to the question and the rationale behind. $\mathcal{R}(s,a)$ is the reward function. Therefore, we have state transition distribution $\mathcal{P}(s_{t}|s_{t-1},a_{t-1})$ and action generation distribution $\mathcal{\pi}(a_{t}|s_{t-1},a_{t-1},q,p_{0},\mathcal{R})$, where $p_{0}$ is the prompt for the Navigator to generate direction $a_{t}$. It is nontrivial to obtain frequent rewards that incentivize self-refinement progress without access to the ground truth. Therefore, we turn to an intrinsically motivated planning algorithm, i.e., Monte-Carlo Tree Search (MCTS) (Kocsis and Szepesvári, 2006; Browne et al., 2012; Swiechowski et al., 2023) to efficiently explore the environment augmenting rewards with auxiliary objectives (Mu et al., 2022; Du et al., 2023b).
@@ 4.2 Multiple-perspective Assessment
Motivated by the empirical results in § 3.1 regarding knowledge-grounding, we propose to employ an advanced consistency-based method as a surrogate for factual correctness when external resources are unavailable. This method considers both intra- and inter-consistency of the generated responses. Specially, we employ the Navigator for $K$ question-oriented direction generation, $a\sim\pi(a_{t}|s_{t-1},a_{t-1},q,p_{0},\mathcal{R})$. These $K$ directions are intended to provide diverse perspectives for problem-solving, with the agreement among guided responses representing inter-consistency. Meanwhile, the confidence in self-consistency (Wang et al., 2023) serves as the measure of intra-consistency.
To integrate consistency considerations into the assessment per reflection iteration, we use intra-consistency to determine whether the Reasoner should accept its initial response. If the intra-consistency for our initial answer surpasses a threshold $T_{0}$, we consider it as the final result; otherwise, we integrate the inter-consistency as an indicator for stopping criteria in subsequent reflection iterations. We derive the final answer when the inter-consistency exceeds $T_{0}$ or when reach the predefined maximum iterations, selecting the final answer with the highest consistency score  [^3] The threshold $T_{0}$ for different models are datasets are set according to the validation performance, details in C.1.. This inter-consistency also becomes part of reward $\mathcal{R}_{\text{consistency}}$ for the current state and contribute to the direction generation. Besides, the intra-consistency value is transformed into verbal form, becoming part of the prompt $p_{0}$ given for Navigator to generate direction. This is inspired by our observation that higher intra-consistency implies a higher likelihood of correctness, so we offer this additional information to assist in feedback generation. This is similar to (Xu et al., 2023), where ICL performance benefits from accessing to the prediction of a supervised fine-tuned smaller model. Our assessment method is different from majority vote, which treats every node with the same weight when aggregating the final result. The comparison results are shown in Table 4.
@@ 4.3 Diverse and Valid Search Space
Obtaining a meaningful and diverse action space is challenging due to the absence of a dense and well-defined reward function in the planning algorithm. One of the predominant reasons is that different action sequences can lead to similar outcomes (Baranes and Oudeyer, 2013). In our context, considering the limitation of LLMs in following instructions, the Reasoner may ignore the differences among multiple directions and generate identical responses merely based on the question. Therefore, some intrinsically motivated reinforcement learning algorithms choose to explore outcomes rather than actions (Oudeyer and Kaplan, 2007; Ladosz et al., 2022). MCTS addresses the limitation of sparse rewards by visiting novel states or transitions through random exploration (Du et al., 2023b). The most popular algorithm in the MCTS family, Upper Confidence Bound for Trees (UCT) (Kocsis and Szepesvári, 2006) is treated as the choice of child node, $\text{UCT}=\overline{\mathcal{R}}_{j}+2C_{p}\sqrt{\frac{2\,\text{In}\,N(n)}{N(n_{j})}}$, where ¯$\overline{\mathcal{R}}_{j}$ is the average reward for child node $j$, while the second term encourages sampling from nodes whose children are less visited. $N(n)$ is the number of times current node (parent) has been visited in previous iterations, and $N(n_{j})$ is times of the child node has been visited. The $C_{p}>0$ is a constant to control balance between exploitation (first term) and exploration (second term). In our case, we specifically promote diversity between the parent and child node, i.e., the response in previous attempt $s_{t-1}$ and the current attempt $s_{t}$. For multiple-choice questions in MMLU, we assess if the predicted choices are the same across two reflection iterations. The discrepancy in responses indicates the alleviation of null direction space and the avoidance of being stuck, especially given the relatively low consistency with the response from the previous iteration. The relationship between task performance and the diversity of responses in the generated tree, as illustrated in Figure 5, confirms our motivation for diversity enhancement.
However, maximizing diversity of outcomes may not always be enough, as less relevant states might be collected (Du et al., 2023b). Therefore, to ensure valid search space, we filter out states whose associated responses are not in the correct form, such as failing to provide a final choice, or refusing to answer questions for moral considerations. For search efficiency, our proposed stopping criteria is to terminate the search process once its inter-consistency surpasses a threshold, thereby avoiding unnecessary search and expansion costs.
@@ 5 Can Mirror Steer LLMs in Iterative Improvements?
We evaluate our proposed Mirror on MMLU and FEVER (Thorne et al., 2018). FEVER is a fact-checking dataset featuring three labels for knowledge-rich statements, i.e., supports, refutes and not enough info. [^4] We also include the evaluation results on GSM8K (Cobbe et al., 2021) in Table A4.
@@ 5.1 Experimental Setup and Results
The evaluation models are GPT-3.5, Llama2-13B-Chat (Touvron et al., 2023), and Vicuna-v1.5-7B (Zheng et al., 2023)  [^5] We denote them as Llama13B and Vicuna13B for simplicity. Experiment details can be found in Appendix C.. We equip the LLMs with different reasoning mechanisms, including Chain-of-Thought (CoT) (Wei et al., 2022), Self-consistency (Wang et al., 2023), Self-Correction (Huang et al., 2023) and Reflexion(w.GT) Shinn et al. (2023). We implement CoT by prompting LLMs to first generate step-by-step thoughts and then generate answers based on those thoughts. We repeat this process for five times, resulting in Self-Consistency(5). The remaining two methods are self-improvement techniques where LLMs are first prompted to generate reflections, followed by updating their current response accordingly if applicable. Self-Correction relies on LLM’s internal knowledge for answer assessment, while Reflexion compares the current answer with the ground truth for evaluation.
Figure 4: Reasoning process of self-correction and Mirror. Text in red are generated directions. Our diversity is characterised in (i) generating directions tailored to questions (ii) encouraging exploration in multiple plausible reasoning paths. The final answer is derived through an agreement among multiple trajectories.
@@
Self-Consistency(5) 0.67 0.68 0.58 0.64 0.61
Self-Correct(2) 0.63 0.62 0.55 0.54 0.55
@@
Self-Consistency(5) 0.45 0.60 0.49 0.57 0.46
Self-Correct(2) 0.42 0.52 0.53 0.45 0.36
@@
Self-Consistency(5) 0.50 0.62 0.53 0.60 0.43
Self-Correct(2) 0.43 0.49 0.42 0.49 0.38
@@ Results.
The results are shown in Table 2. By comparing CoT with Self-Correction, we observe the performance degradation after two rounds of self-Correction across almost all datasets and models. This observation aligns with our findings in §3.1 and in Huang et al. (2023). Equipped with self-consistency(5), significant performance improvements are evident across all settings. Mirror considers additional inter-consistency, achieves the most notable improvements, with a relative increase of more than 15% across the three models. Figure 4 illustrates the reasoning process of Self-correction and Mirror. Both methods fail to answer correctly in the first trial. With question-oriented direction, the Reasoner better identify errors in the initial response, such as, the error in score direction and inconsistency between rationales and selection. The consistency-based criteria built in the tree further improves the fact assessment. During backpropagation, node $s^{(1)}_{1}$ receives a higher reward, leading to the leftmost reasoning path (details of direction $a^{(1)}_{1},a^{(1)}_{2},a^{(2)}_{2}$ and corresponding responses are shown in the text frame). By contrast, Self-correction seems to engage in groundless inference by switching answers without explicit clues. Even comparing Mirror with Relexion(w.GT), we find comparable results for GPT-3.5 on the STEM dataset, for Llama on all datasets except for STEM and for Vicuna on STEM and Humanity. From the perspective of the model, the average improvements over baselines for GPT-3.5 are particularly prominent, partly explained by its better ability to adhere to provided directions. This can also explain the marginal improvements even ground truth are accessible to the smaller models.
@@ 5.2 Analysis
GPT-3.5: CoT 0.68 0.58
+ FixedDirect 0.73 0.60
+ GenerativeDirect 0.78 0.64
Llama13B: CoT 0.46 0.40
+ FixedDirect 0.43 0.39
+ GenerativeDirect 0.49 0.45
Vicuna13B: CoT 0.48 0.42
+ FixedDirect 0.51 0.43
+ GenerativeDirect 0.55 0.45
@@
Figure 5: The Accuracy (acc) and the percentage of samples where the ground truth is included in the tree (ans-presence), with different sizes of search space (Num). Results for GPT-3.5 and Llama13B are in Figure 4(a) and 4(b).
We provide a case to show the effects of diversity in directions (highlighted difference are in bold). The quantitative results of semantic similarity between multiple directions are shown in Table A5.
Case study Question: The price elasticity of demand for a product is greater if - Answer without Direction: C. the number of substitute products is limited. - Answer with Direction: Direction 1: Pay close attention to nuances among options and prioritize selections that resonate with economic principles and the factors influencing consumer behavior. Answer 1: Given the advice, the most suitable choice would be: A. the proportion of the good of the consumer’s budget is high. This aligns with economic principles of demand elasticity. Direction 2: Prioritize options that align with factors impacting consumer responsiveness to price changes, such as budget allocation and substitute availability, while considering the time frame for consumer reactions. Answer 2: Given the advice, the most suitable choice would be: A. the proportion of the good of the consumer’s budget is high. This choice suggests that when a significant portion of a consumer’s budget is allocated to a particular product, they are likely to be more sensitive to price changes.
@@
As discussed in Section 3.1, LLMs struggle to assess the correctness of knowledge-rich statements, a capability that can be consistently enhanced through self-consistency. We further reform the majority-voting assessment process by considering the inter-consistency built in the hierarchical decision-making tree. To study the effects of our answer assessment criteria described in §4.2, we compare them with two other voting methods, i.e., self-consistency and majority vote within our generated tree-trajectories. We average the results from Table 2 for CoT and Self-consistency(5) across four domains in MMLU and denote them as CoT(1) and CoT(5), respectively. We also compare with CoT(15) because our generated trees have at most 3 layers and 5 branches in each layer, so there are 15 candidate nodes. For Majority(tree), we select the final answer through majority-voting among all intermediate nodes in our generated tree-trajectories. The results of different final answer assessments are presented in Table 4. The performance improvements are observed on self-consistency(15) over self-consistency(5), although the improvements percentage isn’t pronounced as that seen in the comparison between self-consistency(5) over CoT. We observe a performance increase after applying majority-voting in the CoT settings, while this simple strategy doesn’t yield improvements in the generated tree. This is because undesirable responses may be generated during the node expanding phase, and majority voting treats all nodes equally. In contrast, our reward-based search tends to focus on reliable nodes with higher confidence in each reflection step, thereby avoiding search cost on less desirable nodes.
@@
GPT-3.5: CoT(1) 0.60 0.58
CoT(5) 0.64 0.61
CoT(15) 0.67 0.62
+Majority(tree) 0.69 0.59
+Reward Search(tree) 0.73 0.64
Llama13B: CoT(1) 0.49 0.40
CoT(5) 0.53 0.46
CoT(15) 0.55 0.48
+Majority(tree) 0.58 0.50
+Reward Search(tree) 0.60 0.54
Vicuna13B: CoT(1) 0.51 0.39
CoT(5) 0.56 0.43
CoT(15) 0.58 0.43
+Majority(tree) 0.59 0.43
+Reward Search(tree) 0.60 0.46
@@
Search Efficiency Analysis.
Mirror is a tree-like searching algorithm that benefits from iterative reasoning process, which enhances the reasoning ability at the computation cost. To mitigate the search cost, we (i) incorporate the Monte-Carlo tree search for its selective search and expansion. (ii) introduce early-stop criteria to encourage a shallow and avoid multiple playouts. We summarise the tree-depth in Table 5. The results below show that our resulting tree, with a maximum depth of 3, is heavily unbalanced and shallow.
STEM Social Hum Others FEVER
Depth=2 0.17 0.12 0.28 0.13 0.04
Depth=3 0.00 0.09 0.20 0.04 0.00
Table 5: Depth of the search tree based on GPT-3.5 across different datasets.